OpenStudy (anonymous):

where does the function intersect the x-axis ? f(x)=(x-1)(x+2)(x-3)

OpenStudy (zzr0ck3r):

we want f(x) = 0 so (x-1)(x+2)(x-3) = 0 this happens when (x-1) = 0 (x+2)=0 (x-3) = 0

OpenStudy (zzr0ck3r):

think about a*b=2 there are infinite things we can make a,b to make this true a = 1/2, b = 4 a = 1/8, b = 16 a = 32, b = 1/16 ...... but when we have ab=0 there are only three things that can happen a=0, b=0, or a and b are both 0

OpenStudy (anonymous):

so for (x-1) i would set x=+1 & for (x+2) i would set x=-2 then for (x-3) i would set x=+3 right ??

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

so f(x)=0*0*0 f(x)=0 so the function intersects with the x-axis 0 times ?